# Adaptive Reduced Tempering For Bayesian inverse problems and rare event simulation

### v1

@@ 2.1 Iterative importance sampling
We consider a vanishing initial inverse temperature $\beta^{(0)}=0$, so that the initial proposal distribution matches the reference distribution $\mu^{(0)}_{\beta^{(0)}}=\pi$. Using this initialization, we will design in Section 2.2 an adaptive procedure which determines at each iteration $k\geq 1$ ($k$ being the number of snapshot evaluations using the true score $S^{\ast}$) some inverse temperature $\beta^{(k)}\in[0,\beta_{\infty}]$ “suited” to the current reduced score $S^{(k)}$. To this temperature is associated the target ${\eta}^{\ast}_{\beta^{(k)}}$ as defined by (1.1) and the proposal distribution $\mu^{(k)}_{\beta^{(k)}}$ as defined by (1.4). Given this adaptive procedure determining the sequence of inverse temperatures $\{\beta^{(k)}\}_{k}$, assume we can compute the normalizations $\{Z^{(k)}_{\beta^{(k)}}\}_{k}$ and that we can draw exactly according to the proposals $X_{k+1}\sim\mu^{(k)}_{\beta^{(k)}}$ in order to compute the sample of snapshots $\left\{(X_{1},S^{\ast}(X_{1})),(X_{2},S^{\ast}(X_{2})),\ldots\right\}$. Then, an iterative importance sampling estimator of the unnormalized target ${\gamma}^{\ast}_{\beta_{\infty}}\displaystyle\mathop{=}^{{\rm def}}Z_{\beta_{\infty}}{\eta}^{\ast}_{\beta_{\infty}}$ is given by
@@
with $H$ denoting the size of the sample used for estimation and where $K_{j_{0}}$ denotes a random iteration index whose definition will be given in due time. The self-normalized estimate of the target ${\eta}^{\ast}_{\beta_{\infty}}$ is
$\hat{\eta}^{H}_{\beta_{\infty}}(\varphi)=\frac{\hat{\gamma}^{H}_{\beta_{\infty}}(\varphi)}{\hat{\gamma}^{H}_{\beta_{\infty}}(\,{\rm l}{\rm l})}.$ (2.2)
Therefore, to enable the evaluation of these estimates, we need at iteration $k$ of the algorithm to i) adapt the inverse temperature $\beta^{(k)}$, so that it is “suited” to the current reduced score function $S^{(k)}$, ii) compute efficiently the normalization $Z_{\beta^{(k)}}$ We discuss these two issues in the next sections.
@@ 2.2 Critical inverse temperature
The idea is to tune at iteration $k$ the inverse temperature $\beta^{(k)}$ with a criterion evaluating that surpassing this inverse temperature should require, in some sense, a better surrogate model, and thus new snapshots. The adjusted proposal $\mu^{(k)}_{\beta^{(k)}}$ will then be used to sample a new snapshot, and obtain an updated score function $S^{(k+1)}$ for the next iteration $k+1$. In short, we are going to develop a heuristic criterion to detect when and where a new snapshot is needed.
@@
The general insight is the following. Consider the following simpler problem: one is given the family of target distributions in the form of (1.1), a given computational budget, and the family of importance sampling proposal distributions in the form of (1.4). One wants to perform some sequential algorithm to sample the distributions $\mu_{\beta}^{(k)}$ along an increasing sequence of inverse temperatures $\beta$ as high as possible, and then use importance sampling. However, we do not want to waste resources for temperature that are “unreasonably” low. As an idealized criterion giving achievable temperatures, we propose the cost of importance sampling of the true target. As is thoroughly argued in [12] (see also [10, 11]), the variance is in general a pessimistic quantification of importance sampling. For a target $\eta$ with proposal $\mu$, it has been proven in  [12] that the cost of importance sampling – in terms of sample size – is roughly given (up to a reasonable tail condition on the log-density $\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu}$), by the exponential of the relative entropy between $\eta$ and $\mu$, i.e., $\mathrm{e}^{\mathrm{Ent}(\eta\mid\mu)},$ where $\mathrm{Ent}(\eta\mid\mu)\displaystyle\mathop{=}^{{\rm def}}\eta(\log\eta)-\eta(\log\mu).$ We will thus call $\mathrm{Ent}(\eta\mid\mu)$ the log cost of importance sampling of the target $\eta$ by the proposal $\mu$. The first non-achievable temperature, hence a kind of critical temperature, is thus given when the latter matches a certain given numerical threshold $c_{1}$. By doing so, we ensure that the snapshot sampled with respect to $\mu_{\beta}^{(k)}$ will not diverge too much from the target distribution $\eta^{\ast}_{\beta}$ at an appropriate inverse temperature $\beta$.
@@
In a practical context, the log cost $\mathrm{Ent}(\eta^{\ast}_{\beta}\mid\mu_{\beta}^{(k)})$ has to be estimated using some prior information on $\eta^{\ast}_{\beta}$, for instance using some form of error quantification. We propose to consider a surrogate of the target $\eta^{\ast}_{\beta}$, denoted by $\check{\eta}_{\beta}^{(k)}$, constructed according to (1.6) with the reduced score $S^{(k)}$ and its error quantification $E^{(k)}$ of the form of (1.2). We stress that this choice of surrogate is merely conventional and other cases could be studied in future work. Our choice is simply a limit case [^2] Taking $S+E$ instead of $S-E$ in the definition of the surrogate $\check{\eta}_{\beta}$ is a possibility but it seems to lead to less robust calculations because it favors states with high errors. We did not studied in detail different options for surrogate targets in the present work. of possible true score imposed by the approximate error bound (1.2). The critical temperature will be estimated using the surrogate target by searching the closest temperature such that the log cost $\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k)}\mid\mu_{\beta}^{(k)}\right)$ matches the threshold $c_{1}$. We emphasize that the irrelevant situation where relative entropy is uniformly zero is excluded by the constraint (1.3), which requires that the error $E^{(k)}$ vanishes on snapshot points $X^{(k)}$ already evaluated at iteration $k$, prohibits constant error quantification functions unless the error estimate $E$ is identically zero. We’d like to point out that this condition is naturally satisfied by many reduced models, and in particular by reduced bases. However, this requirement is not a hard limit on the method, as the choice of the surrogate $\check{\eta}_{\beta}$ can be adapted in other cases.
@@
with the initial temperature at $k=0$ set to zero, i.e., $\beta^{(0)}(c_{1},\beta_{0})\displaystyle\mathop{=}^{{\rm def}}0.$ For convenience, the dependency on $c_{1}$ and $\beta_{0}$ of $\beta^{(k)}$ will be removed if the context makes this clear. $\beta^{(k)}(c_{1},\beta_{0})$ is an inverse temperature such that, in the case where the true target is given by the surrogate $\check{\eta}_{l}^{(k)}$, a log cost of $c_{1}$ is necessary and sufficient in order to perform importance sampling. We will define $\beta_{0}$ later in Section 2.5.
@@ 2.3 Adaptive tempering down to the critical temperature
$\beta_{i+1}:=\max\{\beta:\;\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)\leq c_{2}\,,\,\beta_{i}\leq\beta\leq\beta^{(k)}\},$ (2.5)
where $c_{2}>0$ is some given prescribed level. This one dimensional problem is easily solved since $\beta\mapsto\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)$ is increasing. [^3] Some simple calculation shows that $\partial_{\beta}\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)=(\beta-{\beta_{i}})\textrm{var}_{\ \mu_{\beta}^{(k)}}({S^{(k)}})$ The entropic criterion used in (2.5) may be viewed as an extension to smooth distributions of the criterion used in adaptive multilevel splitting [11]. However, other criteria are possible, see for example [3].
@@ 2.4 Snapshot sampling
As defined Section 2.1, only the snapshots sampled at and after iteration $K_{j_{0}}+1$ are used to build the importance sampling estimator $\hat{\gamma}^{H}_{\beta_{\infty}}$ and $\hat{\eta}^{H}_{\beta_{\infty}}$. The idea here is that the snapshots should be first used to refine the reduced score $S^{(k)}$ around the regime of interest. The snapshots can be used for importance sampling once the reduced score is sufficiently accurate in the sense of a limited variance for importance sampling with distribution $\mu^{(k)}_{\beta^{(k)}}$ of the target $\eta_{\beta^{(k)}}^{\ast}$. We choose to define the trigger index $K_{j_{0}}$ (random index independent of the future snapshots) as the first iteration index $k$ for which among the set of already computed critical inverse temperatures $\{\beta^{(k^{\prime})},\,k^{\prime}\leq k\}$, at least $j_{0}$ of them are equal to the targeted inverse temperature $\beta_{\infty}$, for some prescribed $j_{0}\geq 1$. This snapshot sampling approach corresponds to some extent to a particular setting of the so-called learning function in the active learning literature. [28].
@@
In practice in the SMC implementation, for not too large sample size $N$, we may pick $\tau$ large enough to be virtually equivalent to +$\infty$+ ∞. Then, once the reduced score is related to a temperature close to the target one, new snapshots are sampled according to the importance distribution:
@@ 2.5 Bridging reduced scores
[…]
@@ 2.7 The algorithm
The algorithm presented below gathers the ingredients presented in the previous sections. It is an idealized pseudo-code and, in practice, it is not directly implementable. Indeed, it assumes exact evaluation of expectations and exact sampling according to the proposals. In Section 3, we will propose a practical implementation of this idealized algorithm using a sample of $N$ clones (or particles) SMC methodology. The latter will simulate the flow of proposals for diminishing temperatures using Monte Carlo samples with approximately correct distributions for large sample sizes $N\to+\infty$.
As shown in Appendix A with a martingale argument, the unormalized estimator (2.1) computed by the algorithm below is unbiased
@@
4:    $k\leftarrow k+1$triangleright ( critical temperature and adaptive tempering, see Section 2.2 & 2.3)
@@
$\left\{\begin{array}[]{ll}\beta_{0}=\beta_{\mathrm{bridge}}(k)\\ \beta_{i+1}=\min\{\beta:\;\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)>c_{2}\,,\,\beta_{i}\leq\beta\leq\beta_{\infty}\},\\ \end{array}\right.$ (2.13)
@@
8:    if $j\geq j_{0}$ then draw $X\sim\mu_{\beta^{(k)}}^{(k)},$ else $X\sim\mathrm{e}^{\tau E^{(k)}}\mu_{\beta^{(k)}}^{(k)}/Z^{(k)}_{\tau}.$
9:    Compute the snapshot $S^{\ast}(X)=\mathrm{score}(\Psi(X))$triangleright (stopping criterion, see Section 2.6)
10:
11:    if  $\mathrm{Ent}(\check{\eta}^{(k)}_{\beta^{(k)}}\mid\mu^{(k)}_{\beta^{(k)}})>\epsilon$ or $\beta^{(k)}<\beta_{\infty}$ then
12:         Update reduced model: build $(S^{(k+1)},E^{(k+1)})$ by enriching $(S^{(k)},E^{(k)})$ with $(X,S^{\ast}(X))$.
13:    else
14:         Don’t update the reduced model: set $(S^{(k+1)},E^{(k+1)})\leftarrow(S^{(k)},E^{(k)})$.
15:    end $\triangleright$ ( importance sampling, see Section 2.1)
16:    if  $j>j_{0}$ then
17:         $H\leftarrow H+1$
18:
@@
19:    end if
20:    Set $K\leftarrow K-1$triangleright ( bridging, see Section 2.5)
21:    Find the largest $k^{\prime}\in\llbracket 0,k\rrbracket$ such that exists $\beta^{\prime}\in\mathcal{C}(k^{\prime})$ where:
@@
22:    Simulate the proposal $\mu^{(k+1)}_{\beta_{\mathrm{bridge}}(k+1)}$ starting from $\mu^{(k_{\mathrm{bridge}}(k+1))}_{\beta^{k_{\mathrm{bridge}}(k+1)}}$
23:    Compute using (2.9) the normalization $Z^{(k+1)}_{\beta_{\mathrm{bridge}}(k+1)}$.
24:end while
25:Return $\hat{\gamma}^{H}_{\beta_{\infty}}/\hat{\gamma}^{H}_{\beta_{\infty}}(\,{\rm l}{\rm l})$

### v2

@@ 2.1 Iterative importance sampling
We consider a vanishing initial inverse temperature $\beta^{(0)}=0$, so that the initial proposal distribution matches the reference distribution $\mu^{(0)}_{\beta^{(0)}}=\pi$. Using this initialization, we will design in Section 2.2 an adaptive procedure which determines at each iteration $k\geq 1$ ($k$ being the number of snapshot evaluations using the true score $S^{\ast}$) some inverse temperature $\beta^{(k)}\in[0,\beta_{\infty}]$ “suited” to the current reduced score $S^{(k)}$, in the sense of the log cost of importance sampling defined in (1.5). To this temperature is associated the target ${\eta}^{\ast}_{\beta^{(k)}}$ as defined by (1.1) and the proposal distribution $\mu^{(k)}_{\beta^{(k)}}$ as defined by (1.4). Given this adaptive procedure determining the sequence of inverse temperatures $\{\beta^{(k)}\}_{k}$, consider the idealized case where we can compute the normalizations $\{Z^{(k)}_{\beta^{(k)}}\}_{k}$ and where we can draw exactly according to the proposals $X_{k+1}\sim\mu^{(k)}_{\beta^{(k)}}$ in order to compute the sample of snapshots $\left\{(X_{1},S^{\ast}(X_{1})),(X_{2},S^{\ast}(X_{2})),\ldots\right\}$. Then, an iterative importance sampling estimator of the unnormalized target ${\gamma}^{\ast}_{\beta_{\infty}}\displaystyle\mathop{=}^{{\rm def}}Z_{\beta_{\infty}}{\eta}^{\ast}_{\beta_{\infty}}$ is given by
@@
with $H$ denoting the size of the sample used for estimation and where $K_{j_{0}}$ denotes a random iteration index whose definition will be given in due time. Appendix A shows the consistency of (2.1). The self-normalized estimate of the target ${\eta}^{\ast}_{\beta_{\infty}}$ is
$\hat{\eta}^{H}_{\beta_{\infty}}(\varphi)=\frac{\hat{\gamma}^{H}_{\beta_{\infty}}(\varphi)}{\hat{\gamma}^{H}_{\beta_{\infty}}(\,{\rm l}\hskip-1.99168pt{\rm l})}.$ (2.2)
Therefore, to enable the evaluation of these estimates, we need at iteration $k$ of the algorithm to i) adapt the inverse temperature $\beta^{(k)}$, so that it is “suited” to the current reduced score function $S^{(k)}$ (see Section 2.2), ii) compute efficiently the normalization $Z_{\beta^{(k)}}$ (see Section 2.3).
@@ 2.2 Critical inverse temperature
The idea is to tune at iteration $k$ the inverse temperature $\beta^{(k)}$ with a criterion evaluating that surpassing this inverse temperature should require, in some sense, a better surrogate model, and thus new snapshots. The adjusted proposal $\mu^{(k)}_{\beta^{(k)}}$ will then be used to sample a new snapshot, and obtain an updated score function $S^{(k+1)}$ for the next iteration $k+1$. In short, we are going to develop a heuristic criterion to detect at which tempertaure a new snapshot is needed.
@@
This idea is developed further below. We want to perform some sequential algorithm to sample the distributions $\mu_{\beta}^{(k)}$ along an increasing sequence of inverse temperatures $\beta$ as high as possible, and sometimes possibly use importance sampling to sample according to the target $\eta^{\ast}_{\beta}$ or a variant of it. However, we do not want to waste resources for temperature that are “unreasonably” low because importance sampling becomes unfeasible due to the mismatch between proposal and target. As an idealized criterion giving achievable temperatures, we propose the cost of importance sampling of the true target. As is thoroughly argued in [14] (see also [12, 13]), the variance is in general a pessimistic quantification of importance sampling. For a target $\eta$ with proposal $\mu$, it has been proven in [14] that the cost of importance sampling – in terms of sample size – is roughly given (up to a reasonable tail condition on the log-density [^2] We stress that this tail condition is satisfied by the pair $(\check{\eta}_{\beta},\mu_{\beta})$ we use in practice because the log likelihood is bounded by the finite uniform norm of the error $E$. $\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu}$), by the exponential of the relative entropy between $\eta$ and $\mu$, i.e., $\mathrm{e}^{\mathrm{Ent}(\eta\mid\mu)},$ where we recall that $\mathrm{Ent}(\eta\mid\mu)=\eta(\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu})$. We will thus call $\mathrm{Ent}(\eta\mid\mu)$ the log cost of importance sampling of the target $\eta$ by the proposal $\mu$. The first non-achievable temperature, hence a kind of critical temperature, is thus given when $\mathrm{Ent}(\eta^{\ast}_{\beta}\mid\mu_{\beta}^{(k)})$ matches a certain given numerical threshold $c_{1}$. By doing so, we ensure that the snapshot sampled with respect to $\mu_{\beta}^{(k)}$ will not diverge too much from the target distribution $\eta^{\ast}_{\beta}$ at an appropriate inverse temperature $\beta$.
@@
In a practical context, the log cost $\mathrm{Ent}(\eta^{\ast}_{\beta}\mid\mu_{\beta}^{(k)})$ has to be estimated using some prior information on $\eta^{\ast}_{\beta}$, for instance using some form of error quantification. We propose to consider a surrogate of the target $\eta^{\ast}_{\beta}$, denoted by $\check{\eta}_{\beta}^{(k)}$, constructed according to (1.6) with the reduced score $S^{(k)}$ and its error quantification $E^{(k)}$ of the form of (1.2). We stress that this choice of surrogate is merely conventional and other cases could be studied in future work. Our choice is simply a limit case [^3] Taking $S+E$ instead of SES-E in the definition of the surrogate $\check{\eta}_{\beta}$ is a possibility but it seems to lead to less robust calculations because it favors states with high errors. We did not study in detail different options for surrogate targets in the present work. of possible true score imposed by the approximate error bound (1.2). The critical temperature will be estimated using the surrogate target by searching the closest temperature such that the log cost $\mathrm{Ent}\left(\check{\eta}_{\beta}^{(k)}\mid\mu_{\beta}^{(k)}\right)$ matches the threshold $c_{1}$. We emphasize that the irrelevant situation where relative entropy is uniformly zero is excluded by the constraint (1.3), which requires that the error $E^{(k)}$ vanishes on snapshot points $X^{(k)}$ already evaluated at iteration $k$. This prohibits constant error quantification functions unless the error estimate $E$ is identically zero. We would like to point out that this condition is naturally satisfied by many reduced models, and in particular by reduced bases. However, this requirement is not a hard limit on the method, as the choice of the surrogate $\check{\eta}_{\beta}$ can be adapted in other cases.
@@
with the initial inverse temperature at $k=0$ set to zero, i.e., $\beta^{(0)}(c_{1},\beta_{0})\displaystyle\mathop{=}^{{\rm def}}0.$ Note that the relative entropy in (2.3) is not necessarily monotonic in $\beta$, by definition we simply pick the smallest possible $\beta\geq\beta_{0}$, if it exists. For convenience, the dependency on $c_{1}$ and $\beta_{0}$ of $\beta^{(k)}$ will be removed if the context makes this clear. $\beta^{(k)}(c_{1},\beta_{0})$ is an inverse temperature such that, in the case where the true target is given by the surrogate $\check{\eta}_{l}^{(k)}$, a computational cost of ${\rm e}^{c}_{1}$ is approximately required in order to perform importance sampling. We will define $\beta_{0}$ later in Section 2.5.
@@ 2.3 Adaptive tempering down to the critical temperature
$\beta_{i+1}:=\min\{\beta:\;\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)>c_{2}\,,\,\beta_{i}\leq\beta\leq\beta^{(k)}\},$ (2.5)
where $c_{2}>0$ is some given prescribed level. This one dimensional problem is easily solved since $\beta\mapsto\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)$ is increasing. [^4] Some simple calculation shows that $\partial_{\beta}\mathrm{Ent}\left(\mu_{\beta}^{(k)}\mid\mu_{\beta_{i}}^{(k)}\right)=(\beta-{\beta_{i}})\textrm{var}_{\ \mu_{\beta}^{(k)}}({S^{(k)}})$. The entropic criterion used in (2.5) may be viewed as an extension to smooth distributions of the criterion used in adaptive multilevel splitting [13]. However, other criteria are possible, see for example [5].
@@ 2.4 Snapshot sampling
As defined in Section 2.1, only the snapshots sampled at and after iteration $K_{j_{0}}+1$ are used to build the importance sampling estimator $\hat{\gamma}^{H}_{\beta_{\infty}}$ and $\hat{\eta}^{H}_{\beta_{\infty}}$. The idea here is that the snapshots should be first used to refine the reduced score $S^{(k)}$ around the regime of interest. The snapshots can be used for importance sampling once the reduced score is sufficiently accurate in the sense of a limited variance for importance sampling with distribution $\mu^{(k)}_{\beta^{(k)}}$ of the target $\eta_{\beta^{(k)}}^{\ast}$. We choose to define the trigger index $K_{j_{0}}$ (random index independent of the future snapshots) as the first iteration index $k$ for which among the set of already computed critical inverse temperatures $\{\beta^{(k^{\prime})},\,k^{\prime}\leq k\}$, at least $j_{0}$ of them are equal to the targeted inverse temperature $\beta_{\infty}$, for some prescribed $j_{0}\geq 1$. This snapshot sampling approach corresponds to some extent to a particular setting of the so-called learning function in the active learning literature. [35].
@@
In practice in the SMC implementation, for not too large sample size $N$, we may pick $\tau$ large enough to be equivalent to $\infty$. Then, once the reduced score is related to a temperature close to the target one, new snapshots are sampled according to the importance distribution:
@@ 2.5 Bridging reduced scores
$\text{Find: }\beta\in\mathcal{C}(k_{\mathrm{bridge}})\quad\text{where}\quad k_{\mathrm{bridge}}\displaystyle\mathop{=}^{{\rm def}}\textrm{largest}\,k^{\prime}:\,k^{\prime}\in\llbracket 0,k\rrbracket\text{ s.t. }\mathcal{C}(k^{\prime})\neq\emptyset.$ (2.8)
@@ 2.7 The algorithm
The algorithm presented below gathers the ingredients presented in the previous sections. It is an idealized pseudo-code and, in practice, it is not directly implementable. Indeed, it assumes exact evaluation of expectations and exact sampling according to the proposals. In Section 3, we will propose a practical implementation of this idealized algorithm using a sample of $N$ clones (or particles) in SMC methodology. The latter will simulate the flow of proposals for diminishing temperatures using Monte Carlo samples with approximately correct distributions for large sample sizes $N\to+\infty$.
As shown in Appendix A with a martingale argument, the unnormalized estimator (2.1) computed by the algorithm below is unbiased
@@
4:  $k\leftarrow k+1$ $\triangleright$ ( critical temperature and adaptive tempering, see Section 2.2 & 2.3)
@@
[…]
@@
8:  if $j\geq j_{0}$ then draw $X\sim\mu_{\beta^{(k)}}^{(k)},$ else $X\sim\mathrm{e}^{\tau E^{(k)}}\mu_{\beta^{(k)}}^{(k)}/Z^{(k)}_{\tau}$ end if
9:  Compute the snapshot $S^{\ast}(X)=\mathrm{score}(\Psi(X))$ $\triangleright$ (stopping criterion, see Section 2.6)
10:  if $\mathrm{Ent}(\check{\eta}^{(k)}_{\beta^{(k)}}\mid\mu^{(k)}_{\beta^{(k)}})>\epsilon$ or $\beta^{(k)}<\beta_{\infty}$then
11:  Update reduced model: build $(S^{(k+1)},E^{(k+1)})$ by enriching $(S^{(k)},E^{(k)})$ with $(X,S^{\ast}(X))$
12:   else Don’t update the reduced model: set $(S^{(k+1)},E^{(k+1)})\leftarrow(S^{(k)},E^{(k)})$ end if $\triangleright$ ( importance sampling, see Section 2.1)
13:  if  $j>j_{0}$ then
14:    $H\leftarrow H+1$
15:
@@
16:  end if
17:  Set $K\leftarrow K-1$ $\triangleright$ ( bridging, see Section 2.5)
18:  Find the largest $k^{\prime}\in\llbracket 0,k\rrbracket$ such that exists $\beta^{\prime}\in\mathcal{C}(k^{\prime})$ where:
@@
19:  Simulate the proposal $\mu^{(k+1)}_{\beta_{\mathrm{bridge}}(k+1)}$ starting from $\mu^{(k_{\mathrm{bridge}}(k+1))}_{\beta^{k_{\mathrm{bridge}}(k+1)}}$
20:  Compute using (2.9) the normalization $Z^{(k+1)}_{\beta_{\mathrm{bridge}}(k+1)}$.
21:end while
22:Return $\hat{\gamma}^{H}_{\beta_{\infty}}/\hat{\gamma}^{H}_{\beta_{\infty}}(\,{\rm l}\hskip-1.99168pt{\rm l})$